Valley wood Golf Course offers classes Monday through Saturday. The Saturday classes last 75-minutes while the weekday classes 45-minutes. If they offered 705-minutes of classes last week, what was the maximum number of weekday classes.
a.9
b.14
c.12
d.17

Answer by mducky2(62)


To make the problem easier to solve, we will set it up as the equation of the length of time of each class times the number of classes equals the total amount of minutes. However, since we don't know the number of classes, we'll symbolize our two unknowns with two variables.
75x + 45y = 705
(75x + 45y)/15 = 705/15
5x + 3y = 47
y = (47-5x)/3

It looks like we can't simplify the equation any more, so now it is a matter of trial and error. The minimum number of Saturday classes means the maximum number of weekday classes. We first will test for the maximum by assuming there are no Saturday classes, then will work our way up until x is an integer.

If x = 0
(47-5(0))/3 = 47/3 = 15.6666

If x = 1
(47-5(1))/3 = 42/3 = 14

This works. Therefore, the maximum number of weekday classes is 14, or choice b.